# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v1

@@ 3.1.2. Mean-Teacher Framework
In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017). This framework is characterized by the interplay between a teacher model and a student model. Both networks are initialized with the source-trained model at the beginning of adaptation. The teacher model produces the pseudo-labels for the unlabeled target data, which serve as labels of the unlabeled data to supervise the student model. While the parameters of the student model are optimized via gradient descent, the parameters of the teacher model are updated following an Exponential Moving Average (EMA) strategy based on the student model. Formally, this process can be expressed as follows:
@@
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels generated by the teacher network. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$, respectively. Moreover, the supervision loss $\mathcal{L}_{pl}$, which consists of the rpn loss $\mathcal{L}_{rpn}$ and rcnn loss $\mathcal{L}_{rcnn}$ in faster-rcnn (Ren et al., 2015), is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. Despite the effectiveness of the mean-teacher framework under the static distribution shift, this framework encounters error accumulation and catastrophic forgetting in non-stationary environments. Therefore, we design CTAOD to improve the robustness of feature adaptation in the CTTA setting.
Figure 2. The overview of the proposed CTAOD, which follows the mean-teacher framework. The teacher generates proposals and predictions for the weakly augmented image, whereas the student receives the strongly enhanced images. Our proposed CTAOD includes: (1) The Object-level Contrastive Learning (OCL) module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The dynamic threshold (DT) strategy adjusts category-specific thresholds based on prediction scores; These pseudo-labels are utilized to supervise the student through the threshold. (3) The Data-driven stochastic restoration (DSR) reset the inactive parameter based on the gradient.
@@ 3.2.1. Overview
As indicated in Figure 2, this subsection presents a concise overview of our proposed CTAOD, consisting of Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR). Inspired by UDA methods for object detection (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student model for gradient descent is defined as:
@@
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated in section 3.2.2. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights of the losses. The KL loss is defined as:
@@ 3.2.2. Object-Level Contrastive Learning
In this subsection, we first briefly introduce SimCLR, a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). SimCLR aims to learn high-quality feature representation across differently augmented views of the same image. For the $i^{th}$ image $x_{i}$ in a batch, the CL loss in SimCLR is formulated as follows:
(6) $\mathcal{L}_{cl}(x_{i})=-\log\left(\frac{\exp(sim(z_{i},z_{j})/\tau)}{\sum_{k=1,\ni_{k\neq i}}^{2N}\exp(sim(z_{i},z_{k})/\tau)}\right),$
where $\tau>0$ is a temperature hyper-parameter. $z_{i}$ and $z_{j}$ denote the features of two different augmentations of the same images, serving as the positive pair. At the same time, $z_{k}$ represents one of the features of a different augmented sample in a batch size of $N$. Here, the $sim(\cdot)$ indicates the similarity function, i.e. cosine similarity. Notably, the SimCLR is initially designed for classification tasks, assuming each image pertains to a single category, and requires large batch sizes to ensure sufficient positive and negative pairs for feature representation learning. Consequently, the original SimCLR is not well-suited for object detection, where images typically contain multiple instances, thus requiring significant computational resources to accommodate large batch sizes.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the region proposal network (RPN). Since proposals provide multiple cropped views around the object instance, this strategy does not require a large batch size, thus computationally efficient for online detection model updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. Subsequently, the features are projected into the final object-level features with the same encoder using two fully connected layers. For simplicity, we still denote these two final features as $F_{t}^{T}$ and $F_{t}^{S}$. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
@@
where $l$ indicates the number of the features. This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels. Moreover, the OCL is well integrated into the mean-teacher self-training paradigm as a drop-in enhancement for feature adaptation.
3.2.3. Dynamic Threshold
Due to the presence of domain shifts, pseudo-labels inevitably suffer from label noise. Additionally, the common practice is to use a fixed threshold to remove suspected noisy labels. However, model confidence could differ between different categories and domains, thereby hurting the performance using a fixed threshold. Taking inspiration from the threshold strategy (Wang et al., 2024; Zhao et al., 2023), we design the DT strategy to update the category-specific threshold based on the predicted scores dynamically to minimize noise and select a suitable threshold for each category across different domains in test-time. As shown in Figure 2, the teacher network are first to make the prediction $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images $f_{wek}(x_{t})$ The thresholds for each category are initialized with the same value $\delta_{0}$ at the start of the adaptation, which is updated every iteration by:
(8) $\delta_{t}^{c}\leftarrow\beta\cdot\delta_{t-1}^{c}+(1-\beta)\cdot\epsilon\cdot(\overline{l_{t}^{c}})^{\frac{1}{2}},$
where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l_{t}^{c}}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta$ represents the update rate, and $\epsilon$ provides a linear projection. Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$. For unlabeled data from dynamic environments, this mechanism effectively prevents the threshold from being too high or too low while generating an appropriate threshold for each category.
@@ 3.2.4. Data-Driven Stochastic Restoration
Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022; Zhu et al., 2023) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Taking inspiration from (Wang et al., 2022; Brahma and Rai, 2023), we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for domain at hand.
Specifically, consider a convolution layer of the student model with weight after gradient descent $W_{t-1}$ and its corresponding gradient matrix $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$. First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM (Kirkpatrick et al., 2017):
(9) $F_{t}=G_{t}\odot G_{t}\odot R_{t},$
where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
@@
where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e. $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness, thereby mitigating the forgetting issue in CTTA.
@@ 4. EXPERIMENTS
Table 2. Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the short-term adaptation performance by continually adapting the source-trained model to the twelve corruptions with the largest corruption severity level 5.
Time $t\xrightarrow{\hskip 325.215pt}$ All
Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain
Source (Ren et al., 2015) 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 /
Tent (Wang et al., 2020) 6.8 7.8 7.7 1.3 0.2 6.1 33.1 28.0 2.2 51.1 14.8 11.0 14.2 -0.9
CoTTA (Wang et al., 2022) 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9
SVDP (Yang et al., 2024) 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0
IRG (VS et al., 2023a) 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4
MemCLR (VS et al., 2023b) 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9
Ours 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5
Table 3. Experimental results (mAP@0.5) of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the five corruption ten times with the largest corruption severity level 5. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
Time $t\xrightarrow{\hskip 390.25534pt}$
@@
Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain
Source (Ren et al., 2015) 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 /
Tent (Wang et al., 2020) 35.8 8.1 0.2 29.2 6.2 30.7 10.2 0.8 27.9 12.4 20.4 5.0 0.1 11.5 2.4 12.0 -4.4
CoTTA (Wang et al., 2022) 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7
SVDP (Yang et al., 2024) 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9
IRG (VS et al., 2023a) 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2
MemCLR (VS et al., 2023b) 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6
Ours 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6
In this study, we rigorously evaluate our methodology across four benchmark tasks tailored for CTTA in object detection. Four tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.

### v3

@@ 3.1.2. Mean-Teacher Framework
In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017). Specifically, both networks are first initialized with the source-trained model. The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model. While the parameters of the student model are optimized via gradient descent, the teacher model is updated following an Exponential Moving Average (EMA) strategy based on the student model. Formally, this process can be expressed as follows:
@@
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$. Moreover, the supervision loss $\mathcal{L}_{pl}$ in Faster-RCNN (Ren et al., 2015), which consists of the RPN loss $\mathcal{L}_{rpn}$ and RCNN loss $\mathcal{L}_{rcnn}$, is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments. Therefore, we design CTAOD to fit in the CTTA setting.
@@ 3.2.1. Overview
Figure 2 presents a concise overview of CTAOD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR). Inspired by UDA object detection methods (Li et al., 2022; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student model for gradient descent is defined as:
@@
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated later. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights of the losses. The KL loss is defined as:
@@ 3.2.2. Object-Level Contrastive Learning
SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). SimCLR learns high-quality feature representations from differently augmented views of the same image. Designed for classification tasks, it requires large batch sizes to ensure sufficient positive and negative pairs. Consequently, SimCLR is not well-suited for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detection model updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
@@
where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
3.2.3. Adaptive Monitoring
The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold. The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold.
As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images. Subsequently, We compute the mean scores of all category $\overline{l}_{t}$ in teacher’s predictions and maintain its exponentially moving average $\overline{l}_{t}^{ema}$ through:
(7) $\overline{l}_{t}^{ema}\leftarrow\beta_{s}\cdot\overline{l}_{t-1}^{ema}+(1-\beta_{s})\cdot\overline{l}_{t},$
where $\beta_{s}$ represents the update rate. We introduce a new index $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$. Adaptation resumes when the $\frac{\overline{l}_{t}}{\overline{l}_{t}^{ema}}$ lies within the range $(\frac{1}{\delta_{s}},\delta_{s})$, indicating stability, and pauses when it deviates beyond this range, suggesting changes in data characteristics. This dynamic skipping strategy effectively prevents unstable and noisy adaptation, thus conserving computational resources and mitigating error accumulation.
For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which is updated every iteration by:
(8) $\delta_{t}^{c}\leftarrow\beta_{t}\cdot\delta_{t-1}^{c}+(1-\beta_{t})\cdot\epsilon\cdot(\overline{l}_{t}^{c})^{\frac{1}{2}},$
where $\delta_{t}^{c}$ denotes the threshold of category $c$ at time step t, $\overline{l}_{t}^{c}$ is the mean predicted scores of the category $c$ at time step $t$, $\beta_{t}$ represents the update rate, and $\epsilon$ provides a linear projection. Furthermore, the threshold $\delta_{t}^{c}$ will not change if class $c$ does not exist in prediction $\hat{y}_{t}$, and we set a fixed upper and lower bound $\delta_{max}$ and $\delta_{mini}$. For unlabeled data from dynamic environments, this mechanism effectively generates an appropriate threshold for each category.
@@ 3.2.4. Data-Driven Stochastic Restoration
Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Therefore, we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
Specifically, consider a convolution layer with weight $W_{t-1}$ and its corresponding gradient matrix after gradient descent $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$. First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$, where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
@@
where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e. $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
@@ 4. EXPERIMENTS
Time $t\xrightarrow{\hskip 368.57964pt}$ All
Condition Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast Elastic Pixelate Jpeg Mean Gain Iter.
Source 6.8 8.1 8.0 1.5 0.2 6.8 34.6 30.7 3.0 50.2 17.6 13.5 15.1 / /
CoTTA 7.8 9.0 8.9 1.8 0.3 7.1 38.4 31.1 8.6 49.6 16.2 13.1 16.0 +0.9 6k
SVDP 7.7 10.1 9.7 2.3 0.7 13.0 42.4 45.2 15.4 47.2 21.2 14.8 19.1 +4.0 6k
IRG 8.0 11.0 9.3 3.4 1.2 13.0 37.9 41.3 15.9 38.9 16.9 13.4 17.5 +2.4 6k
MemCLR 8.5 10.4 10.6 2.7 1.1 12.2 41.4 41.6 16.4 43.1 15.4 12.7 18.0 +2.9 6k
Ours-unstop 8.6 12.1 11.7 3.6 1.5 16.7 44.7 48.1 16.7 47.4 22.5 13.9 20.6 +5.5 6k
Ours 8.7 11.7 12.4 3.6 1.5 13.5 43.2 47.3 18.4 47.0 24.8 17.2 20.8 +5.7 4.7k
Table 1. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the performance by continually adapting the source model to the twelve corruptions. “-unstop” denotes no skipping.
Time $t\xrightarrow{\hskip 368.57964pt}$
@@
Condition Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Fog Motion Snow Brightness Defocus Mean Gain Iter.
Source 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 36.1 8.1 0.2 31.0 6.7 16.4 / /
CoTTA 38.2 10.6 0.4 33.8 9.3 40.8 10.4 0.6 36.5 9.2 40.8 10.4 0.5 36.3 9.6 19.1 +2.7 25k
SVDP 36.8 8.8 0.6 43.5 11.2 45.0 15.4 3.8 47.8 19.9 41.5 16.2 4.6 44.6 20.2 25.3 +8.9 25k
IRG 37.9 9.0 0.7 44.3 12.2 45.6 17.7 6.0 46.7 21.9 37.0 16.7 7.1 38.5 20.5 25.6 +9.2 25k
MemCLR 37.7 8.9 0.8 45.5 13.5 45.0 18.1 5.1 46.8 22.4 36.8 18.5 7.5 38.4 21.8 26.0 +9.6 25k
Ours-unstop 39.0 10.4 0.8 48.0 13.8 49.4 18.6 7.7 51.7 25.0 45.7 20.0 11.8 46.4 25.8 29.0 +12.6 25k
Ours 39.2 9.9 0.8 46.7 13.8 49.4 18.5 8.0 52.9 24.7 46.2 19.5 11.3 47.3 29.1 29.2 +12.8 20.1k
Table 2. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the performance by continually adapting the source model to the five corruption ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.